OpenStudy (anonymous):

How much 90 percent vinegar do you have to add to a gallon of 5 percent vinegar to get 20 percent vinegar?

OpenStudy (amistre64):

x(90%) + 1(5%) = (x+1)(20%) .9x + .05 = .2x + .2 .9x -.2x = .2 - .05 .7x = .15 x = .15/.7 = 15/70 = 3/14 gallons 3/14 of a gallon needs to be added. If I did the math in me head correctly :)